# DoFlow: Causal Generative Flows for Interventional and Counterfactual Time-Series Prediction

### v1

@@ Abstract
Time-series forecasting increasingly demands not only accurate observational predictions but also causal forecasting under interventional and counterfactual queries in multivariate systems. We present DoFlow, a flow-based generative model defined over a causal DAG that delivers coherent observational and interventional predictions, as well as counterfactuals through the natural encoding–decoding mechanism of continuous normalizing flows (CNFs). We also provide a supporting counterfactual recovery result under certain assumptions. Beyond forecasting, DoFlow provides explicit likelihoods of future trajectories, enabling principled anomaly detection. Experiments on synthetic datasets with various causal DAG and real-world hydropower and cancer-treatment time series show that DoFlow achieves accurate system-wide observational forecasting, enables causal forecasting over interventional and counterfactual queries, and effectively detects anomalies. This work contributes to the broader goal of unifying causal reasoning and generative modeling for complex dynamical systems.
@@ 1 Introduction
Forecasting time-evolving systems is a foundational task in statistics and machine learning, with decades of progress across classical and modern paradigms. However, in many real-world systems, the goal extends beyond forecasting future observations: we also seek to answer interventional questions—“How will the forecast change under a planned action or modification of certain variables?”—and counterfactual queries—“What would the system have looked like had we intervened differently?” Yet most modern forecasting methods, despite their empirical success, are purely observational: they capture correlations but cannot reason about interventions or alternate scenarios. Classical causal inference methods for time series can model interventions, but they are often restricted to discrete treatments or low-dimensional systems and cannot generate full system trajectories. Bridging these regimes calls for a model that is both causally structured and generative, capable of generating probabilistically consistent trajectories under hypothetical actions.
In this paper, we address this gap by proposing DoFlow, a generative model based on Continuous Normalizing Flows (CNFs) that explicitly embeds a causal directed acyclic graph (DAG) structure. Leveraging the invertibility of CNFs and the temporal conditioning in Neural ODEs, DoFlow provides a unified framework for observational, interventional, and counterfactual forecasting. Beyond forecasting, DoFlow also yields explicit likelihoods for future trajectories, enabling principled change-point and anomaly detection. This framework unifies causal reasoning with generative modeling, advancing toward trustworthy inference and decision support for complex dynamical systems.
@@ 1.1 Related Work
  • Time-series forecasting. Modern approaches can be broadly categorized into four families: (i) classical statistical models such as ARIMA, state-space models, and VAR [69, 23, 26, 73, 59, 63]; (ii) deep sequence models including RNNs/LSTMs and attention-augmented variants [44, 22, 51, 38]; (iii) Transformer-based architectures offering scalability on long sequences [64, 72, 66, 48]; and (iv) recent flow and diffusion models that directly generate trajectories, including autoregressive [53, 54, 20], graph-augmented [14], and horizon-wide designs [1, 60]. These models typically condition on learned context representations and use expressive generative backbones to produce future forecasts, but they remain observational, lacking the ability to simulate interventional and counterfactual trajectories.
@@
  • Causal generative modeling. For static data, interventional and counterfactual queries are traditionally addressed via structural equation models and do-calculus. Recent generative extensions include graph causal encoders [58, 57, 52], diffusion-based causal models [9], and model-agnostic counterfactual generators [30, 16]. However, these methods largely target non-temporal settings and do not capture causal dependencies that evolve over time. By contrast, DoFlow extends causal generative modeling to time-indexed DAGs, enabling coherent interventional and counterfactual forecasting for dynamic systems.
We next discuss three complementary research areas on causal time series—causal effect estimation, counterfactual explainability, and causal discovery—that further contextualize our contribution.
  • Causal effects on time series. Within a causal perspective, prior work has largely focused on modeling treatment effects and unobserved potential outcomes in sequential data, using methods such as Gaussian processes, variational Bayesian models, and classical time-series approaches like ARIMA and Interrupted Time Series [7, 59, 39, 4, 6, 5, 46, 21, 55, 27]. These studies typically examine how binary or discrete, fixed-time actions influence outcomes across a temporal horizon, often by extrapolating an alternative path and contrasting it with the observed post-intervention series. For time-varying treatments, variational approaches estimate individualized effects through balanced representations [36, 42, 34, 67]. Overall, treatment effects are often formalized as $\tau_{t}=\mathbb{E}[Y_{t}|A_{t-1}=j]-\mathbb{E}[Y_{t}|A_{t-1}=k]$, where $Y_{t}$ denotes the outcome and $A_{t}$ the time-varying action.
  • Counterfactual explainability. Beyond treatment effects, another line of work develops counterfactual explanations aimed at interpretability—identifying how past signals must be perturbed to alter a model’s predicted future trajectory or label. For example, “what adjustments to a patient’s breathing signal would lead the model to forecast deeper sleep stages?” Approaches include optimization-based perturbations [3, 17, 25], gradient- or motif-guided objectives [33, 65, 32], and self-interpretable models that couple prediction with counterfactual reasoning [68, 50]. While effective for explaining black-box predictors, these methods do not simulate system-wide trajectories under causal queries.
@@
  • Causal discovery. Inferring causal directed acyclic graphs (DAGs) from observational time series has also been actively explored. Existing approaches include statistical dependence–based tests [45, 8], optimization-based formulations [49], and deep learning–based frameworks [62, 71, 12, 13], which collectively advance the recovery of temporal causal structures under various assumptions.
    Complementary to these efforts, there remains a need for models that explicitly embed causal structure within generative time-series forecasting, thereby enabling interventions on individual variables at arbitrary times and yielding coherent interventional and counterfactual predictions of system-wide trajectories. This direction—though relatively underexplored—is broadly important across domains. In hydropower systems [31], operators may ask how adjusting turbine flow over a given interval will propagate to downstream signals (interventional), or whether an alternative operating schedule could have avoided downstream anomalies (counterfactual). In healthcare [11, 28], dependencies among medication dosage, respiration, heart rate, and oxygenation motivate vital questions such as whether modified dosages during a specific window would have prevented abnormal oxygen levels over time.
@@ 2.1 Settings and goals
We consider a multivariate time series evolving over a causal directed acyclic graph (DAG) with nodes $\{1,\dots,K\}$ in a topologically sorted order. Each node $i$ in the DAG represents a univariate variable $X_{i,t}\in\mathbb{R}$ at discrete time step $t$. Let $\mathbf{X}_{t}:=\{X_{1,t},\dots,X_{K,t}\}$ denote the collection of all nodes at time $t$, and let $X_{\mathrm{pa}(i),t}:=\{X_{j,t}:j\in\mathrm{pa}(i)\}$ denote the values of node $i$’s parents at time $t$. The parent set $\mathrm{pa}(i)$ encodes the structural dependencies that remain fixed across time—for instance, in a physical system, upstream components serve as permanent causal parents of downstream ones. At each step $t$, the value of node $i$ is influenced by its own past trajectory $X_{i,t-}$ and the past trajectories of its parents $X_{\mathrm{pa}(i),t-}$:
$X_{i,t-}:=\{X_{i,s},s<t\},\quad X_{\mathrm{pa}(i),t-}:=\{X_{j,s},j\in\mathrm{pa}(i),s<t\}.$
Formally, we define the underlying structural causal model (SCM) as
@@
where $U_{i,t}$ is an exogenous noise variable independent across nodes and time. Here, the SCM means that $X_{i,t}$ is generated by a function of its parental state $(X_{i,,t-},X_{\mathrm{pa}(i),t-})$ and an unobserved exogenous variable $U_{i,t}$. For training and evaluation, each sequence $\{\mathbf{X}_{1},\dots,\mathbf{X}_{T}\}$ is divided into a context window $\{\mathbf{X}_{1},\dots,\mathbf{X}_{\tau}\}$ used for conditioning, and a forecasting window $\{\mathbf{X}_{\tau+1},\dots,\mathbf{X}_{T}\}$ for prediction.
@@
Our goal is to develop time-series causal prediction under the following two types of regimes:
  • Interventional forecasting. We denote the intervention schedule by $\mathcal{I}\subseteq[K]\times\{\tau+1,\dots,T\}$, where each $(i,t)\in\mathcal{I}$ specifies an intervention on node $i$ at time $t$. Each intervention replaces the system’s natural evolution with a fixed value $\mathrm{do}(X_{i,t}:=\gamma_{i,t})$, collectively written as $\mathrm{do}(X_{\mathcal{I}}:=\gamma_{\mathcal{I}})$. This defines the interventional distribution:
@@ 2.2 Continuous normalizing flow foundations
We propose DoFlow, a time-conditioned continuous normalizing flow (CNF) that autoregressively predicts the future value of each node. At each prediction step $t$, the CNF is conditioned on the past hidden states $H_{t-1}$, which encodes the temporal information up to time t1t{-}1. At each time-series step $t$, a CNF defines a continuous transformation between a base distribution (at s=0s{=}0, typically $\mathcal{N}(0,1)$) and the target distribution (at s=1s{=}1, corresponding to the data at time-series step $t$) through a neural ordinary differential equation (ODE):
$\small\frac{dx_{t}(s)}{ds}=v\!\left(x_{t}(s),s;\,H_{t-1}\right),\quad s\!\in\![0,1],$ (4)
where we learn the velocity field $v$ using a neural network. To predict $x_{t}$, one samples from the base distribution $z=x_{t}(0)\!\sim\!\mathcal{N}(0,1)$ and integrates the above ODE from $s=0$ to $s=1$. We train the velocity field $v$ using Conditional Flow Matching (CFM) loss [40], which directly regresses $v$ onto an analytically defined reference velocity field. Specifically, a reference path $\phi$ interpolates between a base sample $z$ and a data sample $x_{t}$ (typically using linear interpolation), whose derivative $\partial_{s}\phi$ defines the reference velocity. The model minimizes an $L_{2}$ loss between $v$ and $\partial_{s}\phi$, allowing efficient training while preserving the invertible mapping from base distribution to data distribution.
@@ 2.3 Time-conditioned continuous normalizing flow
We learn a continuous normalizing flow (CNF) for each node $i$ to autoregressively generate $X_{i,t}$ for $t=\tau+1,\dots,T$, conditioning each step on the node’s past history and its parents’ past histories. An RNN (can be either an LSTM or GRU [18]) is employed to summarize histories via a recurrent state $h_{i,t}$:
$h_{i,t}=\text{RNN}(\text{concat}\{x_{i,t},c_{i,t}\},h_{i,t-1}),$ (5)
where $c_{i,t}$ denotes optional conditioning variables for the time series, such as embeddings of temporal features. Therefore, for each node $i$ at time $t$, the flow is conditioned on the hidden states:
$H_{i,t-1}:=(h_{i,t-1},h_{\mathrm{pa}(i),t-1}),$ (6)
where $h_{i,t-1}$ and $h_{\mathrm{pa}(i),t-1}$ are the node and parents’ past recurrent states, respectively.
At forecasting time, we use $\hat{H}_{i,t}$ (for interventional forecasting) or $\hat{H}_{i,t}^{\mathrm{CF}}$ (for counterfactual forecasting) to denote the hidden states that are updated from the model predicted values. Besides, $H_{i,t}^{\mathrm{F}}$ denotes the factual hidden state computed from the observed factual trajectory.
Time-conditioned continuous normalizing flow.
Conditioned on histories $H_{i,t-1}$, we train a separate CNF for each node $i$ that are shared across time-series steps $t$. The Neural ODE of the conditional continuous normalizing flow is defined as:
$\small\frac{dx_{i,t}(s)}{ds}=v_{i}(x_{i,t}(s),s;H_{i,t-1}),\quad s\in[0,1],\quad t\in\{\tau+1,\tau+2,\dots,T\},$ (7)
which transforms samples from the base distribution $\mathcal{N}(0,1)$ to the time-series distribution at time $t$. To predict $x_{i,t}$, one samples from the base distribution $z=x_{i,t}(0)\!\sim\!\mathcal{N}(0,1)$ and integrates the ODE from s=0s{=}0 to s=1s{=}1.
@@
At the start of the forecasting window, $H_{i,\tau}$ is initialized from the context sequence $X_{i,1:\tau}$. For later forecasting steps $t\geq\tau+1$, the hidden states $H_{i,t}$ are updated recursively during training using the observed values $(X_{i,t},X_{\mathrm{pa}(i),t})$. The training loss is calculated over the entire forecasting window. For each $t\in\{\tau+1,\dots,T\}$, we define the reference path $\phi$ as a straight-line interpolation between the training sample $x_{i,t}$ and Gaussian noise $z\sim\mathcal{N}(0,I)$:
@@
Here, $\partial_{s}\phi$ represents the reference velocity field, which the model’s learned velocity field $v_{i}$ aims to approximate. Training thus minimizes the squared $L_{2}$ distance between $v_{i}$ and $\partial_{s}\phi$. Therefore, the training loss of the flow using conditional flow matching [40] becomes:
$\small\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{\mathbf{X}_{1:T}\sim p_{\mathcal{X}}}\Bigg[\frac{1}{K(T-\tau)}\sum_{i=1}^{K}\sum_{t=\tau+1}^{T}\;\mathbb{E}_{s\sim\mathcal{U}[0,1],\,z\sim\mathcal{N}(0,I)}\big\|v_{i}\big(\phi(x_{i,t},z;s),\,s;\,H_{i,t-1}\big)-\partial_{s}\phi(x_{i,t},z;s)\big\|_{2}^{2}\Bigg],$ (9)
@@
After training, we obtain a velocity field for each node $i$ on the causal DAG. We next define the forward and reverse processes of the Neural ODE.
@@
We treat the forward process as an encoding operation, denoted by the function $Z_{t}:=\Phi_{\theta}(X_{t};\,H_{t})$. Given the velocity field $v_{i}$, the forward process pushes an observed factual outcome $x_{i,t}^{\mathrm{F}}$ (at $s=0$) to a latent embedding $z_{i,t}^{\mathrm{F}}$ (at $s=1$), conditioned on a factual hidden state $H_{i,t-1}^{\mathrm{F}}$ that summarizes the past observed factual data. Formally,
$z_{i,t}^{\mathrm{F}}=\Phi_{\theta}(x_{i,t}^{\mathrm{F}};\,H_{i,t}^{\mathrm{F}})=x_{i,t}^{\mathrm{F}}+\int_{0}^{1}\!v_{i}\big(x_{i,t}(s),\,s;\,H_{i,t-1}^{\mathrm{F}}\big)\,ds,\ \text{with}\ x_{i,t}(0)=x_{i,t}^{\mathrm{F}}.$ (10)
Through this process, $z_{i,t}^{\mathrm{F}}$ serves as a latent encoding of the factual observation $x_{i,t}^{\mathrm{F}}$.
@@
$\hat{x}_{i,t}=\Phi_{\theta}^{-1}\!\big(z_{i,t};\,\hat{H}_{i,t-1}\big)=z_{i,t}-\!\!\int_{0}^{1}\!v_{i}\big(x_{i,t}(s),\,s;\,\hat{H}_{i,t-1}\big)\,ds,\ \text{with}\ x_{i,t}(1)=z_{i,t}.$ (11)
Here, $\hat{H}_{i,t}$ is autoregressively updated using the predicted values $\hat{x}_{i,t}$ and $\hat{x}_{\mathrm{pa}(i),t}$, and serves as the conditioning state for the next time step.

### v2

# DoFlow: Flow-based Generative Models for Interventional and Counterfactual Forecasting on Time Series

@@ Abstract
Time-series forecasting increasingly demands not only accurate observational predictions but also causal forecasting under interventional and counterfactual queries in multivariate systems. We present DoFlow, a flow-based generative model defined over a causal Directed Acyclic Graph (DAG) that delivers coherent observational and interventional predictions, as well as counterfactuals through the natural encoding–decoding mechanism of continuous normalizing flows (CNFs). We also provide a supporting counterfactual recovery theory under certain assumptions. Beyond forecasting, DoFlow provides explicit likelihoods of future trajectories, enabling principled anomaly detection. Experiments on synthetic datasets with various causal DAG structures and real-world hydropower and cancer-treatment time series show that DoFlow achieves accurate system-wide observational forecasting, enables causal forecasting over interventional and counterfactual queries, and effectively detects anomalies. This work contributes to the broader goal of unifying causal reasoning and generative modeling for complex dynamical systems.
@@ 1 Introduction
Forecasting the evolution of multivariate time series is a central problem in statistics and machine learning, with extensive development across both classical and modern approaches. In the standard setting, a model observes past values and predicts what future values are likely to be under the same conditions that generated the past data. These models are purely observational: they learn correlations from past behavior and extrapolate them into the future.
However, many real-world applications require more than observational forecasting. Practitioners often need to answer “what if” questions about interventions, i.e., how the system would evolve if certain control variables were set differently. In this work, we focus on two types of causal queries:
  • An interventional query asks: “How will the forecast change under a planned modification of certain variables?” Here, the goal is to predict the future trajectory under a chosen sequence of actions. For example, in hydropower time series, turbine-control signals are upstream causes of downstream signals such as power output (Klein et al., 2021). Operators may specify a hypothetical turbine-control plan and ask how all downstream signals will evolve. This contrasts with standard observational forecasters, which, given a fixed past context, produce a fixed forecast. They do not allow one to modify future control plans and inspect how changes propagate through the system.
  • A counterfactual query instead asks: “What would the system trajectories have looked like had we intervened differently?” Here, we have already observed a factual trajectory and ask how the same trajectory would have changed under an alternative intervention. For example, in healthcare (Cheng et al., 2020; Jacob Rodrigues et al., 2020), we may observe a patient’s treatment and outcome trajectory, and then ask whether this particular patient’s outcome trajectory would have been better or worse under a different dosing schedule. Outcomes depend not only on dosing but also on patient-specific factors that are not directly observed (e.g., baseline health), yet these factors are implicitly reflected in the factual trajectory. By conditioning on the factual trajectory, we can infer such unobserved factors and then re-simulate how this same patient’s trajectory would have evolved under the alternative dosing plan.
However, most modern forecasters are observational: they excel at correlational prediction but cannot answer causal queries. Also, to our knowledge, there exists no general framework for counterfactual time-series forecasting yet. Bridging these regimes calls for a model that is both causally structured and generative, capable of generating consistent and system-wide trajectories under causal queries.
Contributions. In this paper, we address this gap by proposing DoFlow, a generative model based on Continuous Normalizing Flows (CNFs) that explicitly embeds a causal directed acyclic graph (DAG) structure. Leveraging the invertibility of CNFs and the temporal conditioning in Neural ODEs, DoFlow provides a unified framework for observational, interventional, and counterfactual forecasting. Beyond forecasting, it also yields explicit likelihoods for future trajectories, enabling principled anomaly detection. By unifying causal reasoning with deep generative modeling, DoFlow advances toward trustworthy inference and decision support in complex dynamical systems.
@@ 1.1 Related Work
  • Time-series forecasting. Modern approaches can be broadly categorized into four families: (i) classical statistical models such as ARIMA, state-space models, and vector autoregression (Young and Shellswell, 1972; Harvey, 1990; Hyndman et al., 2008; Zivot and Wang, 2006; Scott and Varian, 2014; Taylor and Letham, 2018); (ii) deep sequence models including RNNs/LSTMs and attention-augmented variants (Medsker et al., 2001; Graves, 2012; Qin et al., 2017; Lindemann et al., 2021); (iii) Transformer-based architectures offering scalability on long sequences (Vaswani et al., 2017; Zhou et al., 2021; Wu et al., 2021; Nie et al., 2022); and (iv) recent flow and diffusion models that directly generate trajectories, including autoregressive (Rasul et al., 2020; 2021; Gao et al., 2025), graph-augmented (Dai and Chen, 2022), and horizon-wide designs (Alcaraz and Strodthoff, 2022; Shen and Kwok, 2023). These models typically condition on learned context representations and use expressive generative backbones to produce future forecasts, but they remain observational, lacking the ability to simulate interventional and counterfactual trajectories.
@@
  • Causal generative modeling. For static data, interventional and counterfactual queries are traditionally addressed via structural equation models and do-calculus. Recent generative extensions include graph causal encoders (Schölkopf et al., 2021; Sánchez-Martin et al., 2022; Rahman and Kocaoglu, 2024), diffusion and flow-based models (Khemakhem et al., 2021; Javaloy et al., 2023; Chao et al., 2024; Wu et al., 2025), and model-agnostic counterfactual generators (Karimi et al., 2020; Dasgupta et al., 2025). However, these methods largely target non-temporal settings and do not capture causal dependencies over time. A related work studies counterfactual dynamic forecasting (Liu et al., 2023) in constrained physical systems, with interventions only on the initial constrained state. By contrast, DoFlow extends causal generative modeling to general time-indexed DAGs, enabling coherent interventional and counterfactual forecasting across dynamic systems.
We briefly situate DoFlow relative to two complementary lines a practitioner may combine with our method: (i) Causal effect estimation—quantifying how external actions change short-term expected outcomes, and (ii) Causal discovery—recovering a causal DAG from observational data.
  • Causal effects on time series. Prior research on time-series interventions has primarily focused on estimating treatment effects under external actions, using variational Bayesian models, and classical time-series approaches like ARIMA (Brodersen et al., 2015; Scott and Varian, 2014; Linden, 2015; Bernal et al., 2017; Bica et al., 2020b; a; Menchetti et al., 2021; Wijn et al., 2022; Grecov et al., 2022; Giudice et al., 2022; Runge et al., 2023; Hyndman and Rostami-Tabar, 2025). These studies typically model how a discrete action variable $A_{t}\in\{0,\ldots,J{-}1\}$—for example, representing treatment versus control or different policy levels—affects future outcomes of the series. These are often formalized as $\tau_{t}=\mathbb{E}[Y_{t}|A_{t-1}=j]-\mathbb{E}[Y_{t}|A_{t-1}=k]$, where $Y_{t}$ denotes the outcome and $A_{t}$ the time-varying action. Methodologically, modern approaches estimate the expected outcome under alternative actions, often through conditional generative or variational representation-learning models, to compute average or individualized treatment effects (Lim, 2018; Liu et al., 2020; Li et al., 2021; Cao et al., 2023; Wu et al., 2024b; Cinquini et al., 2025).
@@
  • Causal discovery. Inferring causal directed acyclic graphs (DAGs) from observational time series has also been actively explored. Existing approaches include statistical dependence–based tests (Melkas et al., 2021; Bussmann et al., 2021), optimization-based formulations (Pamfil et al., 2020), and deep learning–based frameworks (Tank et al., 2021; Yin and Barucca, 2022; Zhong et al., 2023; Cheng et al., ; Faruque et al., 2024; Cheng et al., 2024), which collectively advance the recovery of temporal causal structures under various assumptions.
However, existing causal effects on time-series methods focus mainly on discrete, fixed-time actions and estimate short-term expected outcome differences. Moreover, to our knowledge, no existing work has modeled the counterfactual trajectory, despite its importance for reliable decision making. Therefore, there remains a need for a generative model that explicitly embeds a causal DAG in time-series, supports interventions on individual continuous variables at arbitrary times, and yields coherent interventional and counterfactual predictions across the system.
@@ 2.1 Settings and goals
We consider a $K$-dimensional multivariate time series evolving over a causal directed acyclic graph (DAG) with nodes $\{1,\dots,K\}$ in a topologically sorted order. We denote by $X_{i,t}\in\mathbb{R}$ the value of node $i$ at time-series step $t$. Let $\mathbf{X}_{t}:=\{X_{1,t},\dots,X_{K,t}\}$ denote the collection of all nodes at time $t$, and let $X_{\mathrm{pa}(i),t}:=\{X_{j,t}:j\in\mathrm{pa}(i)\}$ denote the values of node $i$’s parents at time $t$.
The parent set $\mathrm{pa}(i)$ encodes the structural dependencies that remain fixed—for instance, in a physical system, upstream components serve as permanent causal parents of downstream ones. At each step $t$, the value $X_{i,t}$ depends on its past trajectory $X_{i,t-}$ and the past trajectories of its parents $X_{\mathrm{pa}(i),t-}$:
$X_{i,t-}:=\{X_{i,s},s<t\},\quad X_{\mathrm{pa}(i),t-}:=\{X_{j,s},j\in\mathrm{pa}(i),s<t\}.$ (1)
We assume no within-time-step causal effects, meaning that all causal influences occur with at least one time-step lag. We formalize the dynamics by the structural causal model (SCM):
@@
where $U_{i,t}$ is an exogenous noise independent across nodes and time. Here, $f_{i}$ is a causal mechanism that takes as inputs the histories of $i$ and its parents, plus an exogenous noise.
For training and evaluation, each sequence $\{\mathbf{X}_{1},\dots,\mathbf{X}_{T}\}$ is divided into a context window $\{\mathbf{X}_{1},\dots,\mathbf{X}_{\tau}\}$ used for conditioning and a forecasting window $\{\mathbf{X}_{\tau+1},\dots,\mathbf{X}_{T}\}$ used for prediction. The context window consists of observed past values and is never intervened on. Interventions, if present, are applied within the forecasting window and influence all subsequent steps that we aim to predict. We use uppercase for random variables $\mathbf{X}_{t}$ and lowercase for realizations $\mathbf{x}_{t}$.
@@
Our goal is to develop a framework for causal time-series prediction that answers:
  • Interventional forecasting. We denote the intervention schedule by $\mathcal{I}\subseteq[K]\times\{\tau+1,\dots,T\}$, where each $(i,t)\in\mathcal{I}$ specifies an intervention on node $i$ at time $t$. For each $(i,t)\!\in\!\mathcal{I}$, the variable $X_{i,t}$ is intervened to the value $\gamma_{i,t}$, forming the collection $\gamma_{\mathcal{I}}:=\{\gamma_{i,t}\}_{(i,t)\in\mathcal{I}}$. Each intervention replaces the system’s natural evolution with a fixed value $\mathrm{do}(X_{i,t}:=\gamma_{i,t})$, collectively written as $\mathrm{do}(X_{\mathcal{I}}:=\gamma_{\mathcal{I}})$. This defines the interventional distribution:
@@ 2.2 Continuous normalizing flow foundations
We first introduce the continuous normalizing flow (CNF) foundation of DoFlow. It autoregressively generates future values, with each step conditioned on past temporal information. For each node $i$ in the causal DAG, we learn a time-conditioned CNF that models the conditional distribution of node $i$ at each time step $t$, given its own and its parents’ past temporal context. This context is summarized by a hidden state $H_{i,t-1}$, which aggregates information from the historical sequences $X_{i,t-}$ and $X_{\mathrm{pa}(i),t-}$. In practice, $H_{i,t-1}$ is obtained using RNN-based encoders, as detailed in Section 2.3.
At each time-series step $t$, a CNF defines a continuous transformation between the target distribution (at s=0s{=}0, corresponding to the target data $X_{i,t}$) and a base distribution (at s=1s{=}1, typically $\mathcal{N}(0,1)$) through a neural ordinary differential equation (ODE) (Chen et al., 2018). Here, $s$ denotes the continuous ODE time, which is distinct from the time-series index $t$.
$\frac{dx_{i,t}(s)}{ds}=v_{i}\!\left(x_{i,t}(s),s;\,H_{i,t-1}\right),\quad s\!\in\![0,1],$ (5)
where the velocity field $v_{i}$ is parameterized by a neural network. To predict $x_{i,t}$, we sample $z=x_{i,t}(1)\!\sim\!\mathcal{N}(0,1)$ and integrate the ODE from $s=1$ to $s=0$. We train the velocity field $v_{i}$ using Conditional Flow Matching (CFM) loss (Lipman et al., 2023), which directly regresses $v_{i}$ onto an analytically defined reference velocity field. Specifically, a reference path $\phi$ interpolates between a data sample $x_{i,t}$ and a base sample $z$ (typically using linear interpolation), whose derivative $\partial_{s}\phi$ defines the reference velocity. The model minimizes an $L_{2}$ loss between $v_{i}$ and $\partial_{s}\phi$, allowing efficient training while preserving the invertible mapping from base distribution to data distribution.
@@ 2.3 Time-conditioned continuous normalizing flow
We learn a continuous normalizing flow (CNF) for each node $i$ to autoregressively predict $X_{i,t}$ for $t\in\{\tau+1,\dots,T\}$, conditioning each step on the node’s past history and its parents’ past histories. An RNN (can be either an LSTM or GRU (Dey and Salem, 2017)) is employed to summarize histories of node $i$ via a recurrent state $h_{i,t}$:
$h_{i,t}=\text{RNN}(x_{i,t},h_{i,t-1}).$ (6)
Therefore, for each node $i$ at time $t$, DoFlow is conditioned on the concatenated hidden states:
$H_{i,t-1}:=\mathrm{concat}\big(h_{i,t-1},\,h_{\mathrm{pa}(i),t-1}\big),$ (7)
where $h_{\mathrm{pa}(i),t-1}:=\{h_{j,t-1}:j\in\mathrm{pa}(i)\}$.
At forecasting time, we use $\hat{H}_{i,t}$ (for observational or interventional forecasting) or $\hat{H}_{i,t}^{\mathrm{CF}}$ (for counterfactual forecasting) to denote the hidden states updated from the model predicted values. Importantly, since $\hat{H}_{i,t-1}$ (or $\hat{H}_{i,t-1}^{\mathrm{CF}}$) is autoregressively updated from the model’s generated or intervened values, it carries information about any past interventions applied to node $i$ or its parents. Besides, we denote $H_{i,t}^{\mathrm{F}}$ as the factual hidden state computed from the observed factual trajectory.
Time-conditioned CNF.
We train a separate CNF for each node $i$ that are shared across time-series steps $t$, conditioning each step on $H_{i,t-1}$. The Neural ODE of the conditional CNF is defined as:
$\frac{dx_{i,t}(s)}{ds}=v_{i}(x_{i,t}(s),s;H_{i,t-1}),$ (8)
$s\in[0,1],t\in\{\tau+1,\dots,T\}$, which connects samples from the time-series distribution at time $t$ (at $s=0$) with the base distribution $\mathcal{N}(0,1)$ (at $s=1$).
@@
At the start of the forecasting window, $H_{i,\tau}$ is initialized from the context sequence $(X_{i,1:\tau},X_{\mathrm{pa}(i),1:\tau})$. For later steps $t\geq\tau+1$, the hidden states $H_{i,t}$ are updated autoregressively during training using the observed values $(X_{i,t},X_{\mathrm{pa}(i),t})$. The training loss is calculated over the entire forecasting window. For each $t\in\{\tau+1,\dots,T\}$, we define the reference path $\phi$ as a straight-line interpolation between the training sample $x_{i,t}$ and Gaussian base sample $z\sim\mathcal{N}(0,I)$:
@@
Here, $\partial_{s}\phi$ represents the reference velocity field, which the model’s learned velocity field $v_{i}$ aims to approximate. Training thus minimizes the squared $L_{2}$ distance between $v_{i}$ and $\partial_{s}\phi$. Therefore, the training loss of the flow using conditional flow matching (Lipman et al., 2023) becomes:
$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{\mathbf{X}_{1:T}\sim p_{\mathcal{X}}}$ $\Bigg[\frac{1}{K(T-\tau)}\sum_{i=1}^{K}\sum_{t=\tau+1}^{T}\;$ (10)
$\mathbb{E}_{s\sim\mathcal{U}[0,1],\,z\sim\mathcal{N}(0,I)}\big\|v_{i}\big(\phi(x_{i,t},z;s),\,s;\,H_{i,t-1}\big)-\partial_{s}\phi(x_{i,t},z;s)\big\|_{2}^{2}\Bigg],$
@@
After training, we obtain a velocity field $v_{i}$ for each node $i$ on the causal DAG. We next define the forward and reverse processes of the Neural ODE.
@@
We treat the forward process as an encoding operation, denoted by the function $Z_{t}:=\Phi_{\theta}(X_{t};\,H_{t-1})$. Given the velocity field $v_{i}$, the forward process pushes an observed factual outcome $x_{i,t}^{\mathrm{F}}$ (at $s=0$) to a latent embedding $z_{i,t}^{\mathrm{F}}$ (at $s=1$), conditioned on a factual hidden state $H_{i,t-1}^{\mathrm{F}}$ that summarizes the past observed factual data. Formally,
$z_{i,t}^{\mathrm{F}}=\Phi_{\theta}(x_{i,t}^{\mathrm{F}};\,H_{i,t-1}^{\mathrm{F}})=x_{i,t}^{\mathrm{F}}+\int_{0}^{1}\!v_{i}\big(x_{i,t}(s),\,s;\,H_{i,t-1}^{\mathrm{F}}\big)\,ds,$ (11)
with $x_{i,t}(0)=x_{i,t}^{\mathrm{F}}$. Through this process, $z_{i,t}^{\mathrm{F}}$ encodes the information of the factual observation $x_{i,t}^{\mathrm{F}}$.
@@
$\hat{x}_{i,t}=\Phi_{\theta}^{-1}\!\big(z_{i,t};\,\hat{H}_{i,t-1}\big)=z_{i,t}-\!\!\int_{0}^{1}\!v_{i}\big(x_{i,t}(s),\,s;\,\hat{H}_{i,t-1}\big)\,ds,$ (12)
with $x_{i,t}(1)=z_{i,t}$. Here, $\hat{H}_{i,t}$ is autoregressively updated using the predicted values $\hat{x}_{i,t}$ and $\hat{x}_{\mathrm{pa}(i),t}$, and serves as the conditioning state for the next time step.